OpenStudy (adrianna.gongora):

I need help on this . When solving a radical equation, John and Tim came to two different conclusions. John found a solution, while Tim's solution did not work in the equation. Create and justify two situations: one situation where John is correct and a separate situation where Tim is correct.

OpenStudy (johnweldon1993):

This can happen quite often actually Use the example: \(\large \sqrt{x -1} = x - 7\) If we square both sides we get \(\large x - 1 = x^2 - 14x + 49\) Now we have a quadratic \(\large x^2 - 15x + 50 = 0\) Which we can then solve *quadratic formula, factoring etc...* and find 2 solutions \[\large x = 5\] and \[\large x = 10\] If we plug those both back into our original equation..we find one of them doesn't work

OpenStudy (adrianna.gongora):

Sorry I had to do something.. So would I be able to use this as my answer for the problem?

OpenStudy (adrianna.gongora):

@johnweldon1993